OpenStudy (anonymous):

find two numbers whose sum is 54 and whose difference is 28

OpenStudy (anonymous):

41 and 13

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

nytime...

OpenStudy (anonymous):

Chloe this is how to do it yourself... let the numbers be x and y so ... two numbers whose sum is 54 means x + y = 54 and... two numbers whose difference is 28 meansx + y = 28 if you add the two equations now x + y + x + y = 54 + 28 2x = 82 so x = 41 put this value of x into any of the two equations you made, say x - y = 28 and y is shown to equal 13